OpenStudy (anonymous):

x/5 - 1 = x/5 a x= b solution is all real numbers c no solution

geerky42 (geerky42):

so what do you get?

OpenStudy (anonymous):

i dont really know what to do

OpenStudy (zzr0ck3r):

can anything + 1 be equal to itself?

OpenStudy (anonymous):

no

OpenStudy (zzr0ck3r):

there is the answer:)

OpenStudy (anonymous):

oo

OpenStudy (zzr0ck3r):

the way to see this with the algebra is x/5+ 1 = x/5 subtract x/5 from both sides 1=0 \(\huge YIKES\)

OpenStudy (zzr0ck3r):

change those + to -

OpenStudy (anonymous):

how come ur so smart

OpenStudy (anonymous):

?

geerky42 (geerky42):

Experiences

OpenStudy (zzr0ck3r):

you answered it

OpenStudy (anonymous):

-.-